OpenStudy (anonymous):

In 1980, the U.S. has 5.6% of the world's population. By 1990, the percent had dropped to 5.2%.Assuming the population continues to fall at the same rate, write a linear equation that expresses the U.S. percent of the world's population, y, in terms of the year, x.

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

in 10 yrs it goes down 4%

OpenStudy (astrophysics):

Hey

OpenStudy (anonymous):

hey

OpenStudy (astrophysics):

We can write it in y = mx+b form, which is the equation of a line

OpenStudy (anonymous):

im following

OpenStudy (astrophysics):

So the crucial information which will give us our slope is (5.6%, 1980) and (5.2%, 1990) where the percents can be written as 5.6% = 0.056 and 5.2% = 0.052, we can use this to find the slope, using the formula I showed you a while ago \[m= \frac{ y_2-y_1 }{ x_2-x_1 }\] so here we can use the point slope formula once we find the slope \[y−y_1=m(x−x_1) \]

OpenStudy (anonymous):

then i just plug it in right?

OpenStudy (astrophysics):

Yes find the slope first and then plug it into the point slope formula, once you plug in the point slope formula, set it up as y = mx+b format if you like, depending on your teacher.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (astrophysics):

Np